# PulseReddit: A Novel Reddit Dataset for Benchmarking MAS in High-Frequency Cryptocurrency Trading

### v1

@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated considerable promise in financial decision-making (Liu et al., 2023; Wu et al., 2023). As LLM-based Multi-Agent Systems (MAS) have shown significant potential to enhance the capabilities of single LLMs in addressing complex tasks, some research leverages MAS by assigning distinct roles—such as statistical analysts and trading agents—to different agents, aiming for more detailed and profitable trading strategies (Li et al., 2024a; Yu et al., 2024). Despite these advancements, our review of the literature reveals a significant gap: current MAS predominantly focus on lower-frequency trading strategies, such as daily trading. Consequently, there is a notable scarcity of research applying these multi-agent systems to high-frequency trading (HFT) scenarios, such as 1-hour trading. This research void is critically exacerbated by the absence of publicly available datasets specifically tailored for MAS-based HFT research, particularly those integrating real-time social media signals with market data. This oversight is particularly striking given that HFT is a prevalent strategy actively employed by trading companies to identify and exploit fleeting market opportunities (Aldridge, 2013; Jones, 2013; Nahar et al., 2024).
@@ 4 Related Work
Cryptocurrency Datasets. A growing body of research has contributed public cryptocurrency datasets to advance various tasks such as price forecasting, transaction graph modeling, and account classification (Lin et al., 2020, 2022; Wang et al., 2022; Li et al., 2024a; Huang et al., 2022; Zhang et al., 2024b). Additional efforts address specialized problems like phishing account detection (Chen et al., 2020; Li et al., 2022; Wang et al., 2023). However, most existing datasets are built exclusively on on-chain transaction records, resulting in limited node features and generally lacking high-quality textual or social context. In particular, there is a notable absence of datasets centered on cryptocurrency-related discussions, especially those capturing real-time or short-term market sentiment from social media. Moreover, prior datasets rarely provide the time granularity, off-chain integration, or scale required for rigorous evaluation of high-frequency trading systems. To highlight the novelty of our PulseReddit dataset in addressing these gaps—particularly its focus on Reddit data and its applicability to short-term trading research—we provide a detailed comparison with previous works in Table 10.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated considerable promise in financial decision-making (Liu et al., 2023; Wu et al., 2023). The emergence of LLM-based Multi-Agent Systems (MAS) has significantly advanced complex task solving through role specialization and collaborative frameworks (Wang et al., 2024, 2015), while parallel breakthroughs in LLM reasoning capabilities (Wang et al., 2025a; Chen et al., 2025; Wang et al., 2025b) have enhanced analytical depth. These developments enable researchers to construct sophisticated agent ensembles—assigning distinct roles such as statistical analysts and trading agents—to develop more nuanced trading strategies (Li et al., 2024a; Yu et al., 2024). Despite these advancements, our review of the literature reveals a significant gap: current MAS predominantly focus on lower-frequency trading strategies, such as daily trading. Consequently, there is a notable scarcity of research applying these multi-agent systems to high-frequency trading (HFT) scenarios, such as 1-hour trading. This research void is critically exacerbated by the absence of publicly available datasets specifically tailored for MAS-based HFT research, particularly those integrating real-time social media signals with market data. This oversight is particularly striking given that HFT is a prevalent strategy actively employed by trading companies to identify and exploit fleeting market opportunities (Aldridge, 2013; Jones, 2013; Nahar et al., 2024).
@@ 4 Related Work
Cryptocurrency Datasets. A growing body of research has contributed public cryptocurrency datasets to advance various tasks such as price forecasting, transaction graph modeling, and account classification (Lin et al., 2020, 2022; Wang et al., 2022; Li et al., 2024a; Huang et al., 2022; Zhang et al., 2024b; Luo et al., 2024). Additional efforts address specialized problems like phishing account detection (Chen et al., 2020; Li et al., 2022; Wang et al., 2023). However, most existing datasets are built exclusively on on-chain transaction records, resulting in limited node features and generally lacking high-quality textual or social context. In particular, there is a notable absence of datasets centered on cryptocurrency-related discussions, especially those capturing real-time or short-term market sentiment from social media. Moreover, prior datasets rarely provide the time granularity, off-chain integration, or scale required for rigorous evaluation of high-frequency trading systems. To highlight the novelty of our PulseReddit dataset in addressing these gaps—particularly its focus on Reddit data and its applicability to short-term trading research—we provide a detailed comparison with previous works in Table 10.
@@ 5 Conclusion
Acknowledgements. This research was funded by JST SPRING, Japan, Grant Number JPMJSP2180.